2. Form the pair of linear equations in the following problems, and find their solutions (if they exist) by the elimination method:

(i) If we add 1 to the numerator and subtract 1 from the denominator, a fraction reduces to 1. It becomes if we only add 1 to the denominator. What is the fraction?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: Let the fraction be a/b According to the given information, (a+1)/(b-1) = 1 => a – b = -2 ………………………………..(i) a/(b+1) = 1/2 => 2a-b = 1…………………………………(ii) When equation (i) is subtracted from equation (ii) we get, a = 3 …………………………………………………..(iii) When a = 3 is substituted in equation (i) we get, 3 – b = -2 -b = -5 b = 5